OpenStudy (anonymous):

solve the equation: 5x^3=5x(x+12)

OpenStudy (anonymous):

the farthest i got was up to this step 5x^2+ 60x-5x^3

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

Well aperntly x=0 is a solution if not devide by x and you get a quadratic equation.

OpenStudy (anonymous):

5x^2+60x-5x^3=0

OpenStudy (anonymous):

5x^3=5x(x+12) = 5x^2 = 5x + 60 x^2 - x - 12 =0

OpenStudy (anonymous):

ok so from 5x^3=5x(x+12) then you move the 5x^3 so you get 5x(x+12)-5x^3=0 so now you distribute the 5x that gives you 5x^2 +60x-5x^3=0

OpenStudy (anonymous):

I believe this answer is going to have 3 solutions.

OpenStudy (anonymous):

right because of the x^3

OpenStudy (anonymous):

-5x^2 +5x +60 = 0 to find the last two roots.

OpenStudy (anonymous):

factorable.

OpenStudy (anonymous):

Do you understand how to do that now @Mberdeja501 ?

OpenStudy (anonymous):

what where did the - come from? @Algebraic! ?

OpenStudy (anonymous):

from x(-5x^2 +5x +60) =0

OpenStudy (anonymous):

When I worked it I came out with solutions of 0,-3,4. Is that what you got also?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great!!

OpenStudy (anonymous):

factorable how? im lost :( haha @Algebraic!

OpenStudy (anonymous):

(-5x +20)(x+3)

OpenStudy (aravindg):

have you got the answer or need further assistance?

OpenStudy (anonymous):

Let me know what I need to go over again.. I will be more than glad to. :)

OpenStudy (anonymous):

oops ok i get it now. thanks

OpenStudy (anonymous):

No problem @Mberdeja501!! Glad I could help you!!